# Classifier-free graph diffusion for molecular property targeting

### v1

@@ Abstract
This work focuses on the task of property targeting: that is, generating molecules conditioned on target chemical properties to expedite candidate screening for novel drug and materials development. DiGress is a recent diffusion model for molecular graphs whose distinctive feature is allowing property targeting through classifier-based (CB) guidance. While CB guidance may work to generate molecular-like graphs, we hint at the fact that its assumptions apply poorly to the chemical domain. Based on this insight we propose a classifier-free DiGress (FreeGress), which works by directly injecting the conditioning information into the training process. CF guidance is convenient given its less stringent assumptions and since it does not require to train an auxiliary property regressor, thus halving the number of trainable parameters in the model. We empirically show that our model yields up to 79% improvement in Mean Absolute Error with respect to DiGress on property targeting tasks on QM9 and ZINC-250k benchmarks. As an additional contribution, we propose a simple yet powerful approach to improve chemical validity of generated samples, based on the observation that certain chemical properties such as molecular weight correlate with the number of atoms in molecules.
@@ 1 Introduction
The problem of generating molecules with desired chemical properties is crucial to enable fast candidate screening in the early stages of drug (Dara et al. 2021) or novel materials (Liu et al. 2017) development. Although the advent of deep learning has achieved remarkable accomplishments in computational chemistry, this strategic endeavor remains an area of active research. In the past, much of the researchers’ focus has been on the task of property optimization, which is based on modifying an already generated molecule such that it acquires some solicited chemical property (Jin, Barzilay, and Jaakkola 2018). However, driven by the success of conditional generative models (Radford et al. 2021), an equally important – though slightly different – task has emerged recently: property targeting, that is, to natively generate molecules which satisfy pre-specified chemical desiderata (Sousa et al. 2021).
Among the plethora of deep generative models that have been developed in the last decade, some are based on textual representation of molecules such as SMILES (Weininger 1988) or SELFIES (Krenn et al. 2022), while others are based on generating the molecular graph directly. Regarding the latter category, which is of interest for this work, previous approaches have employed autoregressive generation conditioned on latent representations (e.g. Gómez-Bombarelli et al. 2018), as well as one-shot generation of the adjacency and node feature matrices of the molecular graph (e.g. De Cao and Kipf 2022).
@@
Denoising Diffusion Probabilistic Models (DDPMs) (Ho, Jain, and Abbeel 2020) have been shown to achieve state-of-the-art performance in conditional generation, for example to synthesize images from user-provided textual guidance (Saharia et al. 2022). Due to their flexibility and excellent performance, DDPMs have been extended to the chemical domain and applied to with textual and graph-based approaches to tasks such as distribution learning, property optimization, and property targeting with promising results (Runcie and Mey 2023).
@@
In particular, DiGress (Vignac et al. 2023) is one of the first successful applications of DDPMs for molecular generation. Under the hood, DiGress is based on a discrete diffusion process (Austin et al. 2021) which gradually applies noise to a molecular graph according to a transition matrix, while denoising is performed by a graph transformer network (Dwivedi and Bresson 2020). The most interesting feature of DiGress is the possibility to perform conditional generation for property targeting through classifier-based (CB) guidance (Dhariwal and Nichol 2021). Loosely speaking, CB guidance requires to train a separate classifier to predict the conditional information from noisy samples, and to inject the resulting gradients back into the reverse process to bias the generative process. While successful to some extent, CB guidance has been shown to be inherently limited by a) the necessity of training a separate property predictor, which defies the purpose of having a single conditional generative model in the first place, and b) the fact that the gradients of the property predictor are not always informative and may lead the generative process astray. Due to these limitations, in the current DDPMs practice classifier-free (CF) guidance is often preferred. The idea behind CF guidance is to directly incorporate the conditioning vector as input to train the conditional DDPM. With respect to CB guidance, CF guidance has been shown to enable more stable training and better generative performance in general (Ho and Salimans 2022).
@@
With DiGress, and molecular generation specifically, CB guidance is further limited by the fact that the auxiliary model is a regressor whose predictions (the chemical properties) are assumed to be normally distributed even for noisy graphs which are chemically invalid. This motivates our first contribution, which consists of the development and implementation of CF guidance for DiGress, called FreeGress. Experimentally, we show that switching from CB to CF guidance is beneficial to improve at property targeting. In particular, we evaluated FreeGress against DiGress on the QM9 (Ramakrishnan et al. 2014) and ZINC-250k (Gómez-Bombarelli et al. 2018) datasets, where we queried the models to generate molecules with properties as close as possible to a target specification. Comparing the mean absolute error between the target properties and the properties of the generated molecules, FreeGress significantly outperformed DiGress, with improvements up to $79\%$ in the most favourable case. Besides improving performances, FreeGress does not require an auxiliary property regressor, which cuts the number of trainable parameters by half. Furthermore, guided by the observation that certain chemical properties relate to the molecular graph size (the number of atoms in the molecule), we also propose to improve the generative process by first learning the probability of sampling a certain number of nodes given the target property, and then using samples from this distribution to constrain the size of the graph to be generated. Through experiments, we show that this simple method allows to generate more chemically valid graphs without sacrificing performance.
@@ 2 Background and related works
2.1 Notation
For the purposes of this work, a molecular graph with $n$ atoms is a tuple $G=(\bm{X},\bm{\textsf{E}})$ where $\bm{X}\in\mathbb{R}^{n\times a}$ is a node matrix whose $i$-th row, indicated as $\bm{x}_{i}\in\mathbb{R}^{a}$, is the one-hot encoding of the atomic type of the $i$-th atom. Possible atomic types are specified by the set $\mathcal{X}$ with cardinality $a$. Similarly, $\bm{\textsf{E}}\in\mathbb{R}^{n\times n\times b}$ is an adjacency tensor (also called edge tensor) that jointly represents the molecular connectivity and the type of bond occurring between each pair of atoms. In practice, the $(i,j)$-th entry of E, indicated as $\bm{e}_{ij}\in\mathbb{R}^{b}$, contains the one-hot encoding of the bond type occurring between the $i$-th and $j$-th atoms. Possible bond types are specified by the set $\mathcal{E}$ with cardinality $b$, where the absence of a bond is itself considered a bond type. In the following we will use the term “node” interchangeably with “atom”, and “edge” interchangeably with “bond”. Conditioning vectors will be generally represented as vectors $\bm{y}\in\mathbb{R}^{d}$, where $d$ is usually a small integer (1 or 2 in our experiments).
@@ 2.2 Deep generative models for molecules
Historically, the first approach to use a deep generative model to produce unseen molecules has been proposed by Gómez-Bombarelli et al. (2018). Essentially, the model is a VAE where both the encoder and decoder are recurrent neural networks trained to reconstruct SMILES strings. In particular, the decoder is essentially an autoregressive SMILES language model capable of generating a molecule by sampling one SMILES character at a time. Although simplistic, this model still remains one of the most efficient. The model was designed to allow property optimization through Bayesian optimization in latent space. Over the years, several variants to address limitations such as low validity rates (Kusner, Paige, and Hernández-Lobato 2017) or the inability to handle more complex chemical substructures at the vocabulary level (Podda, Bacciu, and Micheli 2020) have been developed. Other methods related to this generative flavour use the alternative language of SELFIES instead, which is deemed to be more robust than SMILES as every possible character sequence defines a valid molecule. A different branch of methods is focused on generating the molecular graph directly. In this case, two possibilities have been explored: generating the graph autoregressively – either atom-by-atom (Li et al. 2018), edge-by-edge (You et al. 2018b; Goyal, Jain, and Ranu 2020; Bacciu and Podda 2021), or motif-by-motif (Jin, Barzilay, and Jaakkola 2020) – or directly predicting the adjacency and node feature matrices (De Cao and Kipf 2022). On graph-based models, property optimization has been performed with simple hill climbing in latent space (guided by a property predictor), up to iterative reinforcement learning approaches which assign higher rewards to chemically appealing molecules (You et al. 2018a). In particular, Jin, Barzilay, and Jaakkola (2020) propose to cast the property optimization problem as a molecular translation problem, where the difference between two molecules (one with suitable chemical properties, the other chemically similar but with poor chemical properties) is mapped to a Gaussian manifold. Latent vectors sampled from this prior can be used to modify unsuitable molecules into suitable ones. Lately, DDPMs that generate the molecular graph have started to be used for distribution learning as well as property optimization tasks. The primary application of DDPMs is 3D structure generation, which is useful to handle a different class of tasks such as ligand generation conditioned on the coordinates of a protein pocket (Baillif et al. 2023). Among DDPM-based models, DiGress was the first to be applied to 2D molecule generation, as well as one of the first to formalize the property targeting task.
@@ 2.3 DiGress: Denoising Diffusion for Graphs
Here we briefly describe DiGress (Vignac et al. 2023), a discrete DDPM for graphs which is the starting point of this study. Preliminarily, we provide a brief introduction to DDPM to set the notation and concepts for understanding our contribution.
Denoising Diffusion Probabilistic Models.
DDPMs consist of an untrained forward process $q(\bm{x}^{t}|\bm{x}^{t-1})$ paired with a parameterized reverse process $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$. The former iteratively perturbs the initial data point $\bm{x}^{0}$ to transform it into an $\bm{x}^{T}$ which is akin to Gaussian noise, while the second is trained to incrementally remove the noise from $\bm{x}^{T}$ until $\bm{x}^{0}$ is restored. Data generation from a trained DDPM is performed by sampling noise $\bm{\tilde{x}}^{T}\thicksim\mathcal{N}(0,\bm{I})$, where $\mathcal{N}$ is an isotropic Gaussian, and iteratively applying the denoising model for $T$ steps until a new datapoint $\bm{\tilde{x}}^{0}\thicksim q$ is obtained.
Conditioned generation in DDPMs is achieved by injecting a guidance vector, or guide $\bm{y}$, to obtain a conditioned reverse process $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})$. CB guidance (Dhariwal and Nichol 2021) refactors the reverse process as $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})p_{\phi}(\bm{y}|\bm{x}^{t-1})$, where $p_{\phi}(\bm{y}|\bm{x}^{t-1})$ is an auxiliary classifier trained to predict the guide from a noisy version of the input. A scaled version of the gradient $\nabla_{\bm{x}^{t}}$ of $g_{\phi}$ is then added to the denoising gradient in order to magnify the conditioning signal. CB guidance is limited by the fact that it requires to train the auxiliary classifier, without the possibility of exploiting pre-trained models; moreover, since only few parts of the noisy $\bm{x}^{t}$ are actually useful to predict $\bm{y}$, the resulting gradient could yield undesirable directions in input space. In contrast, CF guidance (Ho and Salimans 2022) works by jointly optimizing $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})$, the conditioned model, and $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$, the unconditioned model, at the same time. During sampling, the reverse process is computed as $(1+w)\,p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})-w\,p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$, which is a barycentric combination between the conditioned and unconditioned predictions with weight $w\in\mathbb{R}_{+}$.
DDPMs defined as above are not suited for discrete data such as graphs, since discrete distributions cannot be effectively corrupted using Gaussian noise. Instead, discrete DDPMs (Austin et al. 2021), including DiGress, inject noise in the data at time-step $t-1$ by multiplying it against a transition matrix $\bm{Q}^{t}$, where $[\bm{Q}^{t}]_{ij}=q(x^{t}=j|x^{t-1}=i)$, which loosely speaking specifies the probability of changing $\bm{x}$ from category $i$ to category $j$.
DiGress Forward process.
DiGress applies noise to a molecular graph $G$ at time-step $t$ according to the following process:
$q(G^{t}|G^{t-1})=(\bm{X}^{t-1}\bm{Q}^{t}_{\bm{X}},\bm{\textsf{E}}^{t-1}\bm{Q}^{t}_{\bm{\textsf{E}}}),$ (1)
@@
where $\bm{Q}^{t}_{X}$ and $\bm{Q}^{t}_{\bm{\textsf{E}}}$ are transition matrices applied to the node matrix and the edge tensor, respectively. In particular, $\bm{Q}^{t}_{X}$ is chosen from the following distribution:
@@
DiGress Reverse process.
DiGress reconstructs a molecular graph from a noisy sample with the following reverse process:
$p_{\theta}(G^{t-1}|G^{t}),$ (4)
which can be thought of as the joint distribution of the individual nodes and edges given the current graph. By assuming independence, we can rewrite the joint as a product of conditionals:
$\prod_{1\leq i\leq n}p_{\theta}(\bm{x}_{i}^{t-1}|G^{t})\prod_{1\leq i,j\leq n}p_{\theta}(\bm{e}_{ij}^{t-1}|G^{t}).$ (5)
In turn, each node conditional is computed by marginalizing over the node types (resp. edge types for the edge conditionals). Specifically, we have:
$p_{\theta}(\bm{x}_{i}^{t-1}|G^{t})=\sum_{\bm{x}\in\mathcal{X}}q(\bm{x}_{i}^{t-1}|\bm{x}_{i}^{t},\bm{x}^{0}_{i}=\bm{x})\,f_{\theta}(\bm{x}^{0}_{i}=\bm{x}|G^{t}),$ (6)
where $f_{\theta}$ is computed by a neural network trained to infer the node types of the true graph $G^{0}$ from a noisy version $G^{t}$. The edges marginal is computed similarly, by predicting the true edge types from their noisy intermediates. CB guidance for DiGress is achieved by specifying the conditioned reverse process as:
$p_{\theta}(G^{t-1}|G^{t})p_{\phi}(\bm{y}|G^{t-1}),$ (7)
where $p_{\phi}(\bm{y}|G^{t-1})\propto\exp(-\lambda\langle\triangledown_{G^{t}}||\bm{y}-g_{\phi}(G^{t})||^{2},G^{t-1}\rangle)$, $g_{\phi}$ is an auxiliary property regressor, and $\lambda$ is a hyper-parameter that magnifies the conditioning gradient.
Limitations.
While the development of unconditioned DiGress appears intuitive, the use of CB guidance for the conditional variant leads to design choices and assumptions that are unrealistic in the chemical domain. Firstly, CB guidance relies on an external predictor that learns chemical properties from noisy graphs, which implies that chemically invalid molecules are unnaturally related to valid ones by being assigned the same properties. Secondly, the distribution learned by the auxiliary property regressor is assumed to be Normal with mean $p(\bm{y}|G^{t})$, approximated by $g_{\phi}$ by minimizing a mean squared error objective. This assumption is unsupported by premise for most denoising steps (where $G^{t}$ usually represents an invalid molecule), since it implies a distribution of chemical properties for an object that does not even exist in chemical space. Moreover, even if two molecules with graphs $G^{t}$ and $G^{t-1}$ are both valid and similar, their chemical properties can be drastically different, since molecular property landscapes are in general non-smooth (Aldeghi et al. 2022), violating again the normality assumption. This discussion motivates our intention to develop CF guidance for DiGress with the objective of making the conditional process simpler and at the same time appropriate for the chemical context where it is applied.
3 Classifier-Free Graph Diffusion Process
In this section, we first work out how to incorporate CF guidance into the general architecture of DiGress. Then, we detail the specifics of the neural architecture implementing CF guidance. Finally, we describe a chemically-motivated variant of our approach based on predicting the number of nodes of the graph to be generated from the conditioning vector.

### v2

# Matteo Ninniri,Marco Podda,Davide Bacciu

@@ Abstract
This work focuses on the task of property targeting: that is, generating molecules conditioned on target chemical properties to expedite candidate screening for novel drug and materials development. DiGress is a recent diffusion model for molecular graphs whose distinctive feature is allowing property targeting through classifier-based (CB) guidance. While CB guidance may work to generate molecular-like graphs, we hint at the fact that its assumptions apply poorly to the chemical domain. Based on this insight we propose a classifier-free DiGress (FreeGress), which works by directly injecting the conditioning information into the training process. CF guidance is convenient given its less stringent assumptions and since it does not require to train an auxiliary property regressor, thus halving the number of trainable parameters in the model. We empirically show that our model yields significant improvement in Mean Absolute Error with respect to DiGress on property targeting tasks on QM9 and ZINC-250k benchmarks. As an additional contribution, we propose a simple yet powerful approach to improve the chemical validity of generated samples, based on the observation that certain chemical properties such as molecular weight correlate with the number of atoms in molecules.
Keywords:
Graph Neural Networks Generative Models Deep Learning.
@@ 1 Introduction
00footnotetext: This preprint has not undergone peer review (when applicable) or any post-submission improvements or corrections. The Version of Record of this contribution is published "in Machine Learning and Knowledge Discovery in Databases. Research Track", and is available online at https://doi.org/10.1007/978-3-031-70359-1_19
@@
Generating molecules with desired chemical properties is crucial to enable fast candidate screening in the early stages of drug [4] or novel materials [14] development. Although the advent of deep learning has achieved remarkable accomplishments in computational chemistry, this strategic endeavor remains an area of active research. In the past, much of the researchers’ focus has been on the property optimization task, which is based on modifying a generated molecule such that it acquires some solicited chemical property [11]. However, driven by the success of conditional generative models [16], an equally important – though slightly different – task has emerged recently: property targeting, that is, to natively generate molecules which satisfy pre-specified chemical desiderata [21].
@@
Denoising Diffusion Probabilistic Models (DDPMs) [9] have been shown to achieve state-of-the-art performance in conditional generation, for example to synthesize images from user-provided textual guidance [20]. Due to their flexibility and excellent performance, DDPMs have been extended to the chemical domain, being applied to tasks such as distribution learning, property optimization, and property targeting with promising results [19].
@@
DiGress [24] is one of the first successful applications of DDPMs to molecular generation. Under the hood, DiGress is based on a discrete diffusion process [2] which gradually applies noise to a molecular graph according to a transition matrix, while denoising is performed by a graph transformer network [6]. The most interesting feature of DiGress is the possibility to perform conditional generation for property targeting through classifier-based (CB) guidance [5]. Loosely speaking, CB guidance requires to train a separate classifier to predict the conditional information from noisy samples, and to inject the resulting gradients back into the reverse process to bias the generative process. While successful to some extent, CB guidance has been shown to be inherently limited by i) the necessity of training a separate property predictor, which defies the purpose of having a single conditional generative model in the first place, and ii) the fact that the gradients of the property predictor are not always informative and may lead the generative process astray. Due to these limitations, classifier-free (CF) guidance for DDPMs is often preferred. The idea behind CF guidance is to directly incorporate the conditioning vector as input to train the conditional DDPM. With respect to CB guidance, CF guidance has been shown to enable more stable training and better generative performance in general [10].
@@
With DiGress for molecular generation specifically, CB guidance is further limited by the fact that the auxiliary model is a regressor whose predictions (the chemical properties) are assumed to be normally distributed even for noisy graphs which are chemically invalid. This motivates our first contribution, which consists of the development and implementation of CF guidance for DiGress, called FreeGress. Experimentally, we show that switching from CB to CF guidance is beneficial to improve at property targeting. In particular, we evaluated FreeGress against DiGress on the QM9 [17] and ZINC-250k [8] datasets, where we queried the models to generate molecules with properties as close as possible to a target specification. Comparing the mean absolute error between the target properties and the properties of the generated molecules, FreeGress significantly outperformed DiGress, with improvements up to $88\%$ in the most favourable case. Besides improving performance, FreeGress does not require an auxiliary property regressor, reducing the number of trainable parameters. Furthermore, guided by the observation that certain chemical properties relate to the molecular graph size (the number of atoms in the molecule), we also propose to improve the generative process by first learning the probability of sampling a certain number of nodes given the target property, and then using samples from this distribution to constrain the size of the graph to be generated. Through experiments, we show that this simple method allows to generate more chemically valid graphs without sacrificing performance. Our code and supplementary material are available at https://github.com/Asduffo/FreeGress.
@@ 2 Background and related works
2.0.1 Notation.
We represent a molecular graph with $n$ atoms as a tuple $G=(\bm{X},\bm{\textsf{E}})$ where $\bm{X}\in\mathbb{R}^{n\times a}$ is a node matrix whose $i$-th row, indicated as $\bm{x}_{i}\in\mathbb{R}^{a}$, is the one-hot encoding of the atomic type of the $i$-th atom. Possible atomic types are specified by the set $\mathcal{X}$ with cardinality $a$. Similarly, $\bm{\textsf{E}}\in\mathbb{R}^{n\times n\times b}$ is an adjacency tensor (also called edge tensor) that jointly represents the molecular connectivity and the type of bond occurring between each pair of atoms. The $(i,j)$-th entry of E, indicated as $\bm{e}_{ij}\in\mathbb{R}^{b}$, contains the one-hot encoding of the bond type occurring between the $i$-th and $j$-th atoms. Possible bond types are specified by the set $\mathcal{E}$ with cardinality $b$, where the absence of a bond is itself considered a bond type. In the following we will use the term “node” interchangeably with “atom”, and “edge” interchangeably with “bond”. Conditioning vectors will be generally represented as vectors $\bm{y}\in\mathbb{R}^{d}$, where $d$ is usually a small integer.
@@ 2.1 Deep generative models for molecules
The first approach to use a deep generative model to produce unseen molecules has been proposed by [8]. Essentially, the model is a VAE where both the encoder and decoder are recurrent neural networks trained to reconstruct SMILES strings. Other methods related to this generative flavour use the alternative language of SELFIES instead, which is deemed to be more robust than SMILES as every possible character sequence defines a valid molecule. A different branch of methods is focused on generating the molecular graph directly. On graph-based models, property optimization has been performed with simple hill climbing in latent space (guided by a property predictor), up to iterative reinforcement learning approaches which assign higher rewards to chemically appealing molecules [25]. Lately, DDPMs that generate the molecular graph have started to be used for distribution learning as well as property optimization tasks.
2.2 Denoising Diffusion Probabilistic Models
DDPMs consist of an untrained forward process $q(\bm{x}^{t}|\bm{x}^{t-1})$ and a parameterized reverse process $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$. The former iteratively perturbs the initial data point $\bm{x}^{0}$ to transform it into an $\bm{x}^{T}$ akin to Gaussian noise, while the second is trained to incrementally remove the noise from $\bm{x}^{T}$ until $\bm{x}^{0}$ is restored. Generation from a trained DDPM amounts to sample $\bm{\tilde{x}}^{T}\thicksim\mathcal{N}(0,\bm{I})$ from an isotropic Gaussian $\mathcal{N}$ and iteratively applying the denoising model for $T$ steps until a new sample $\bm{\tilde{x}}^{0}\thicksim q$ is obtained. Conditioned generation in DDPMs is achieved by injecting a guidance vector, or guide $\bm{y}$, to obtain a conditioned reverse process $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})$. DDPMs defined as above are not suited for graphs, since Gaussian noising does not work effectively on discrete distributions.
2.2.1 Classifier-based guidance
CB guidance [5] refactors the reverse process as $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})p_{\phi}(\bm{y}|\bm{x}^{t-1})$, where $p_{\phi}(\bm{y}|\bm{x}^{t-1})$ employs an auxiliary classifier trained to predict the guide from a noisy version of the input. A scaled version of the regressor gradient $\nabla_{\bm{x}^{t}}$ is used to magnify the conditioning signal. CB guidance is limited by the fact that it requires to train the auxiliary classifier, without the possibility of exploiting pre-trained models; moreover, since only few parts of the noisy $\bm{x}^{t}$ are actually useful to predict $\bm{y}$, the resulting gradient could yield undesirable directions in input space.
2.2.2 Classifier-free guidance
Differently from the CB approach, CF guidance [10] jointly optimizes $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})$, the conditioned model, and $p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$, the unconditioned model, at the same time. During sampling, the reverse process is computed as the barycentric combination $(1+w)\,p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t},\bm{y})-w\,p_{\theta}(\bm{x}^{t-1}|\bm{x}^{t})$ between the conditioned and unconditioned predictions with weight $w\in\mathbb{R}_{+}$.
@@ 2.3 DiGress: Denoising Diffusion for Graphs
Here we briefly describe DiGress [24], a discrete DDPM for graphs which is the starting point of this study. In the forward process, DiGress injects noise in the data at time-step $t-1$ by multiplying it with a transition matrix, which loosely speaking specifies the probability of transitioning from one node type to another. The forward process is applied until the training graph is completely corrupted. The reverse process implemented by DiGress is specified as $p_{\theta}(G^{t-1}|G^{t})$, which is factored (by assuming node and edge independence) as a product of conditionals of the individual nodes and edges given the current graph. In turn, each node conditional is obtained by marginalizing over the node types (resp. edge types for the edge conditionals). Specifically, the nodes (resp. edge) marginal is computed by predicting the true node (resp. edge) types from their noisy intermediates. For conditioned generation, DiGress employs CB guidance. Specifically, the conditioned reverse process is formulated as:
$p_{\theta}(G^{t-1}|G^{t})p_{\phi}(\bm{y}|G^{t-1}),$ (1)
where $p_{\phi}(\bm{y}|G^{t-1})\propto\exp(-\lambda\langle\triangledown_{G^{t}}||\bm{y}-g_{\phi}(G^{t})||^{2},G^{t-1}\rangle)$, $g_{\phi}$ is a property regressor, and $\lambda$ is a hyper-parameter that magnifies the conditional gradient.
2.3.1 Limitations
While the development of unconditioned DiGress appears intuitive, the use of CB guidance for the conditional variant leads to design choices and assumptions that are unrealistic in the chemical domain. Firstly, CB guidance relies on an external predictor that learns chemical properties from noisy graphs, which implies that chemically invalid molecules are unnaturally related to valid ones by being assigned the same properties. Secondly, the distribution learned by the auxiliary property regressor is assumed to be Normal with mean $p(\bm{y}|G^{t})$, approximated by $g_{\phi}$ by minimizing a mean squared error objective. This assumption is unsupported by premise for most denoising steps (where $G^{t}$ usually represents an invalid molecule), since it implies a distribution of chemical properties for an object that does not even exist in chemical space. Moreover, even if two molecules with graphs $G^{t}$ and $G^{t-1}$ are both valid and similar, their chemical properties can be drastically different, since molecular property landscapes are in general non-smooth [1], violating again the normality assumption. This discussion motivates our intention to develop CF guidance for DiGress with the objective of making the conditional process simpler and at the same time appropriate for the chemical context where it is applied.
[…]
Figure 1: A depiction of FreeGress. The forward process, which gradually corrupts a molecule into a random graph, goes from left to the right. The reverse process, which denoises the original graph, goes from right to left. Note that the reverse process allows for a conditioning vector $\bm{y}$ and a number of nodes $n$ sampled from a trained neural network $p_{\varepsilon}$.
3 Classifier-Free Graph Diffusion
This section presents our first contribution, a classifier-free DDPM for conditioned molecular generation named FreeGress. A high-level summarization of the proposed model is sketched in Figure 1.
3.0.1 Forward process
The forward process of FreeGress is analogous to that of DiGress, and essentially consists of applying $T$ noising steps to an input molecular graph $G$ according to the following process:
$q(G^{t}|G^{t-1})=(\bm{X}^{t-1}\bm{Q}^{t}_{\bm{X}},\bm{\textsf{E}}^{t-1}\bm{Q}^{t}_{\bm{\textsf{E}}}),\;1\leq t\leq T$ (2)
@@ 3.0.1 Forward process
where $\bm{Q}^{t}_{X}$ and $\bm{Q}^{t}_{\bm{\textsf{E}}}$ are transition matrices applied to the node matrix and the edge tensor, respectively. In particular, $\bm{Q}^{t}_{X}$ is formulated as:
@@
3.0.2 Reverse Process
The key impact of CF guidance is instead on the adaptive reverse process. To this end, we begin by rewriting the conditioned reverse process as:
$p_{\theta}(G^{t-1}|G^{t},\bm{y}),$ (5)
such that it now incorporates the guide $\bm{y}$ as input. Then, we assume independence between nodes and edges and factorize as the following product of conditionals:
$p_{\theta}(G^{t-1}|G^{t},\bm{y})=\prod_{1\leq i\leq n}p_{\theta}(\bm{x}_{i}^{t-1}|G^{t},\bm{y})\prod_{1\leq i,j\leq n}p_{\theta}(\bm{e}_{ij}^{t-1}|G^{t},\bm{y}).$ (6)
Following [7], we rewrite the nodes marginal as:
$p_{\theta}(\bm{x}_{i}^{t-1}|G^{t},\bm{y})=\sum_{\bm{x}\in\mathcal{X}}q(\bm{x}_{i}^{t-1}|\bm{x}_{i}^{t},\bm{x}_{i}^{0}=\bm{x})\,f_{\theta}(\bm{x}^{0}_{i}=\bm{x}|G^{t},\bm{y}),$ (7)
where $f_{\theta}$ is a neural network that predicts the true node types $\bm{x}^{0}$ using both the current noisy graph $G^{t}$ and the guide $\bm{y}$. Similarly, the edges marginal is:
$p_{\theta}(\bm{e}_{ij}^{t-1}|G^{t},\bm{y})=\sum_{\bm{e}\in\mathcal{E}}q(\bm{e}_{ij}^{t-1}|\bm{e}_{ij}^{t},\bm{e}_{ij}^{0}=\bm{e})\,f_{\theta}(\bm{e}_{ij}^{0}=\bm{e}|G^{t},\bm{y}).$ (8)
At this point, we have, in principle, a classifier-free DDPM for graphs. However, conditioning might not work as expected in this scenario, since the model could learn to completely ignore the guide $\bm{y}$. This behaviour has been observed, e.g., by [22]. To prevent this, we leverage conditional dropout [10], where the conditioning vector $\bm{y}$ is replaced, with probability $\rho$, by a non-parametrized placeholder $\bm{\bar{y}}$. To summarize, the final vector $\bm{y}$ employed in the equations (and, consequently, in the neural network) is chosen as follows:
$\bm{y}\leftarrow\begin{cases}\bm{\bar{y}}&\text{with probability }\rho\\ \bm{y}&\text{with probability }1-\rho\end{cases}$ (9)
Intuitively, the placeholder signals the network to perform inference without a guidance vector, as if the process was unconditioned. As a result, the model becomes capable of performing both conditioned and unconditioned sampling. To complete the derivation of FreeGress, we build on [10] and express $f_{\theta}$ as:
$p_{\theta}(\bm{x}^{0}_{i}|G^{t})+s(p_{\theta}(\bm{x}^{0}_{i}|G^{t},\bm{y})-p_{\theta}(\bm{x}^{0}_{i}|G^{t}))$ (10)
which is computable since the model can output the unconditioned probability using the placeholder in place of the guide or, more formally, since ${p(\bm{x}^{0}_{i}|G^{t})}\approx{p_{\theta}(\bm{x}^{0}_{i}|G^{t},\bm{\bar{y}}})$. The process is symmetric for the edges $\bm{e}\in\mathcal{E}$. Finally, to add more flexibility, we follow [22] and we parametrize $\bm{\bar{y}}$ so that it is trainable.
The major difference between [10] and our work is the fact that the former is designed for continuous signals, such as images, and their neural network is trained to predict the Gaussian noise $\epsilon$ corrupting the input signal $\bm{x}^{t}$ given both the latter and the conditioning vector $\bm{y}$. FreeGress, instead, is designed for discrete graphs and its neural network is trained to infer $G^{0}$ given the corrupted graph $G^{t}$ and conditioning vector $\bm{y}$. An alternative formula, proposed by [22], is shown in Equation 11:
$\log{p(\bm{x}^{0}_{i}|G^{t})}+s(\log{p(\bm{x}^{0}_{i}|G^{t},\bm{y})}-\log{p(\bm{x}^{0}_{i}|G^{t})}).$ (11)
Note how setting the hyper-parameter $s$ to zero is equal to disabling the conditional dropout altogether.